What volume of a 0.251 M perchloric acid solution is required to neutralize 14.6 mL of a 0.123 M barium hydroxide solution?

mL perchloric acid
Answer #1

Balanced chemical reaction is:

2 HClO4 + Ba(OH)2 → 2 H2O + Ba(ClO4)2

moles of NBa(OH)2 = Molarity x Volume in Liter = (0.123 M) x (0.0146 L) = 0.0018 mol

moles of HClO4 needed = (2 /1) x (0.0018 mol) = 0.0036 mol

Given; Molarity of HClO4 = 0.251 M

Volume of HClO4 = moles / Molarity = (0.0036 mol) / (0.251 M) = 0.0143 Liter = 14.3 mL